# A Weight-aware-based Multi-source Unsupervised Domain Adaptation Method for Human Motion Intention Recognition

### v1

@@ Abstract
Accurate recognition of human motion intention (HMI) is beneficial for exoskeleton robots to improve the wearing comfort level and achieve natural human-robot interaction. A classifier trained on labeled source subjects (domains) performs poorly on unlabeled target subject since the difference in individual motor characteristics. The unsupervised domain adaptation (UDA) method has become an effective way to this problem. However, the labeled data are collected from multiple source subjects that might be different not only from the target subject but also from each other. The current UDA methods for HMI recognition ignore the difference between the each source subject, which reduces the classification accuracy. Therefore, this paper considers the differences between source subjects and develops a novel theory and algorithm for UDA to recognize HMI, where the margin disparity discrepancy (MDD) is extended to multi-source UDA theory and a novel weight-aware-based multi-source UDA algorithm (WMDD) is proposed. The source domain weight, which can be adjusted adaptively by the MDD between each source subject and target subject, is incorporated into UDA to measure the differences between source subjects. The developed multi-source UDA theory is theoretical and the generalization error on target subject is guaranteed. The theory can be transformed into an optimization problem for UDA, successfully bridging the gap between theory and algorithm. Moreover, a lightweight network is employed to guarantee the real-time of classification and the adversarial learning between feature generator and ensemble classifiers is utilized to further improve the generalization ability. The extensive experiments verify theoretical analysis and show that WMDD outperforms previous UDA methods on HMI recognition tasks.
@@ I Introduction
The model-free methods map the sensing signals into target HMI directly without parameter identification [13, 14, 15]. The above model-free methods can achieve good performance on specific subject but fail to perform well on cross-subjects. Due to the difference in individual motor characteristics, such as kinematic properties, most current methods require labeling a large amount of data and training specific classifiers for each individual, which is burdensome [9]. Unsupervised domain adaptation (UDA) is a common method to solve the above problem [16, 17, 18, 19]. In this scenario, the source subjects (domains) data are labeled, and the target subject (domain) data are unlabeled. However, there is a discrepancy between the source and target domains. To alleviate the performance reduction caused by this discrepancy, single-source UDA (SUDA) utilizes the adversarial-based methods to align two domains [20, 21, 22] or explores different metric learning schemes to minimize this divergence [23, 24, 25].
In actual HMI recognition tasks, the labeled data are collected from multiple source subjects that might be different not only from the target subject but also from each other. However, the above UDA methods for HMI recognition don’t consider the difference between the source subjects, which may result in a sub-optimal solution to the problem and hinder the improvement of classifier performance [26]. Therefore, this paper aims to propose a multi-source UDA (MUDA) method, which considers the difference between the source subjects, for HMI recognition to further improve the classifying accuracy in the target subject. In MUDA field, there are following two key challenges that need to be addressed.
@@
Figure 1: The overview of the weighted multi-source unsupervised domain adaptation (WMDD) method for human motion intention recognition. The training data are from $N$ source subjects with labels and one target subject without labels. The features of the sensor signal are extracted by feature generator $\phi$. The final goal is to use feature generator and ensemble classifiers to classify the target subject intention accurately. Firstly, the $j$-th auxiliary classifier $\bm{f^{\prime}}_{j}$ optimized by maximizing the discrepancy $\mathcal{P}$. Secondly, the $N$ source domain weight $\bm{\alpha}$ is computed according to the estimated margin disparity discrepancy, and then the feature generator $\phi$ and $K$ classifiers $\bm{f}$ are trained to minimize total loss $\mathcal{L}_{\bm{f},\phi}$. Then, motivated by adversarial learning, the classifiers and generator are optimized by minimizing and maximizing classifiers discrepancy.
@@
For challenge 2, the current MUDA methods are mainly constructed based on the distribution-weighted hypothesis, where the source domain is represented by a convex combination of multi-source domains [30]. Most scholars [31, 32, 33, 34] calculated the source domain weight through the estimated discrepancy between source and target domains and optimized the weighted sum loss of multiple sources to improve the accuracy of classification. However, the weight learned by the above method lacks interpretability and cannot capture the complex relationship between domains [35], and there is a gap between the divergence in theories and algorithms for MUDA methods.
To solve the above problems, we extend the margin disparity discrepancy in SUDA theory to MUDA and propose a novel weighted MUDA algorithm (WMDD) motivated by the margin disparity discrepancy [23] and distribution-weighted hypothesis [30], where the new generalization bound of MUDA is proven and the new method for determining the source domain weight is given. Moreover, this paper theoretically and empirically validates that WMDD can learn diverse features and adapt well to unknown target subject in HMI recognition tasks. Fig. 1 shows the overview of the proposed WMDD. The contributions of this paper are given below.
@@
    A theory of MUDA is developed based on margin disparity discrepancy and a novel weight-aware-based MUDA algorithm is proposed for HMI recognition.
@@
    A lightweight network is designed to guarantee the real-time of classification and adversarial learning is utilized to further improve the generalization ability.
@@
II Related Works
This section presents a brief overview of the literature in the area of human motion intention (HMI) recognition and multi-source unsupervised domain adaptation (MUDA).
@@ II-A Human Motion Intention Recognition
Model-free methods based on deep networks, which can obtain higher-level features from sensor signals without domain-specific knowledge, have achieved great success in recent years, such as CNN [36], LSTM [37], CNN-BiLSTM [38]. However, the above methods are difficult to perform well on cross-subjects due to individual differences. Domain adaption is an effective method to cross-domain (subject) problems. Zhang et al.[18] incorporated an unsupervised cross-subject adaptation method to predict the HMI of the target subject without labels. Zhang et al.[16] proposed a novel non-adversarial cross-subject adaptation method for HMI recognition. The ensemble diverse hypotheses method was designed to mitigate the cross-subject divergence [17]. However, the above methods ignore the difference between each source subject, essentially a single-source UDA method, which may cause a sub-optimal solution to the problem. Few multi-source UDA studies have focused on recognizing HMI. Therefore, this paper aims to propose a novel multi-source UDA method for HMI recognition. To the best of our knowledge, WMDD is the first work about multi-source UDA used for HMI recognition.
II-B Multi-source Unsupervised Domain Adaptation
@@
Domain Adaptation Theory: Ben-David et al.[39, 40] conducted the pioneering theoretical works in domain adaptation field. They used the $\mathcal{H}\Delta\mathcal{H}$ divergence to replace the traditional distribution discrepancies and overcame the difficulties in estimation from finite samples. Mansour et al.[41] extended the zero-one loss of [39] to the general loss function of binary classification and developed a generalization theory. Kuroki et al.[42] proposed a more tractable source-guided discrepancy by fixing the hypothesis of [41] to the ideal source minimizer. Zhang et al.[23] extended the theory of [40] to multiple classes by introducing margin disparity discrepancy (MDD), which can characterize the difference of the multi-class scoring hypothesis. Zhang et al.[24] proposed multi-class scoring disagreement divergence based on MDD [23], that can characterize element-wise disagreements of multi-class scoring hypotheses by aggregating violations of absolute margin.
Multi-source Unsupervised Domain Adaptation Algorithm: Existing MUDA algorithms focused on aligning the distribution of each pair of source and target to reduce their domain shift by minimizing the combined discrepancy between source and target domains. Wen et al.[43] proved that the target loss is bounded by the weighted source loss and weighted discrepancy distance, and optimized networks by minimizing source loss and discrepancy. Yao et al.[44] quantified the importance of different source domains and aligned the source and target distributions by minimizing maximum mean discrepancy. Liu et al.[45] integrated multi-source domains into a single domain and aligned this distribution with the target distribution. Chen et al.[46] optimized the networks through minimizing the source loss and the Hellinger distance between source and target domains. Chen et al.[47] optimized the mixing weights and the network by minimizing the source mixture loss and the Pearson divergence.
Our work is different from the above ones in several key aspects. 1) The source domain weight is adaptively adjusted through the margin disparity discrepancy estimated by auxiliary classifiers instead of optimized constantly by solving unconstrained problem [47], which reduces the computation cost and enhances algorithm stability. 2) The theory and algorithm of WMDD are based on each single source distribution instead of the whole multi-source class-conditionals [44] or the whole multi-source joint distributions [46]. WMDD fully considers the differences between each source subject, which is beneficial for improving classification accuracy. 3) The distribution disparity metric is margin disparity discrepancy that contains more generalization bound information, rather than maximum mean discrepancy [44], the Hellinger distance [46], or Pearson divergence [47]. 4) A lightweight network is employed to guarantee the real-time of classification, instead of using multiple complex networks [43, 45] that is difficult to guarantee the real-time.
@@ III-A Problem Formulation
Fig. 2 compares single-source unsupervised domain adaptation (SUDA) and multi-source unsupervised domain adaptation (MUDA). In MUDA, we consider $\mathcal{C}$ class, $N$ source domains $\{\mathcal{D}_{sj}\}_{j=1}^{N}$ and one target domain $\mathcal{D}_{t}$ problem. There are two different but related distributions over $\mathcal{X}\times\mathcal{Y}$, namely the $j$-th source $P_{j}$ and target $Q$, where $\mathcal{X}$ denotes input space and $\mathcal{Y}$ denotes output space, which is $\{0,1\}$ in binary classification. The learner is trained on labeled data $\mathcal{D}_{sj}=\{(\bm{x}_{i}^{sj},y_{i}^{sj})\}_{i=1}^{|\mathcal{D}_{sj}|}=(\bm{X}_{sj},\bm{Y}_{sj})$ sampled from source distribution $P_{j}$ and unlabeled data $\mathcal{D}_{t}=\{\bm{x}_{i}^{t}\}_{i=1}^{|\mathcal{D}_{t}|}=\bm{X}_{t}$ sampled from target distribution $Q$. The goal of MUDA is to learn the hypothesis space $\mathcal{H}$ of labeling function $h:\mathcal{X}\rightarrow\mathcal{Y}$ to minimize expected target error
@@
where $L$ is the loss function. Ben-David et al.[40] used zero one loss of the form $\mathbb{I}[h(\bm{x})\neq y]$ to represent $L$, where $\mathbb{I}$ is the indicator function. Following [24], we consider the hypothesis space $\mathcal{F}$ that contains scoring function $\bm{f}:\mathcal{X}\rightarrow\mathbb{R}^{|\mathcal{Y}|}=\mathbb{R}^{\mathcal{C}}$, where the outputs denote the prediction confidence. Then the labeling function can be induced
@@
where $f_{y}$ denotes the $y$-th component of vector function $\bm{f}$. Zhang et al.[23] defined the margin of a hypothesis $\bm{f}$ at a labeled sample $(x,y)$ as $\omega_{\bm{f}}(\bm{x},y)=[f_{y}(\bm{x})-\max_{y^{\prime}\neq y}f_{y^{\prime}}(\bm{x})]/2$. Then, the margin loss of $\bm{f}$ can be denoted as
@@
Figure 2: The comparison of single-source unsupervised domain adaptation (SUDA) and multi-source unsupervised domain adaptation (MUDA). The source and target distributions of SUDA are not matched well. For MUDA, the target distribution hardly matches all subject distributions, and the discrepancy between the target distribution and each subject distribution might be different. Different shapes represent different data categories.
Margin disparity discrepancy (MDD) [23], which contains more generalization information, is used to measure the distribution divergence between $P$ and $Q$, defined as
@@
The proof for Theorem 1 can be found in Appendix -A. In the following section, we give the cross-domain generalization bound for MUDA based on Theorem 1.
@@ III-B Generalization Bound for MUDA
In MUDA, some source domains, which are more relevant to target domain than the others, might be more important to classify. Therefore, we use domain weight $\alpha_{j}\geq 0$ to describe this and give following theorem.
@@
The proof of Theorem 2 can be found in Appendix -B. Given the fixed $\beta$, the target error (generalization error) $\mathcal{L}_{Q}\left(h_{\bm{f}}\right)$ is determined by the distribution discrepancy $d^{\mu}_{\bm{f}}(P_{j},Q)$ between $P_{j}$ and $Q$ and the expected loss $\mathcal{L}_{{P}_{j}}^{\mu}(\bm{f})$ over the $j$-th source domain. In Eq. (7), the term $d^{\mu}_{\bm{f}}(P_{j},Q)$ bounds the performance gap caused by domain shift. The smaller $\mathcal{L}_{{P}_{j}}^{\mu}(\bm{f})$ indicates better performance of scoring function $\bm{f}$ on the $j$-th source domain. Note that, the above theorem only considers the ideal and complete distribution. However, the collected data fails to cover the entire distribution space in actual learning.
Therefore, we further consider the sampling error and give the empirical form of Theorem 2. First, Rademacher complexity that is widely used in the generalization theory is introduced to measure the richness of a certain hypothesis space [48, 23]. The definition is given below:
@@
$\Omega_{1}(\mathcal{F})=\Big{\{}(\bm{x},y)\rightarrow\bm{f}_{y}(\bm{x})~{}|~{}\bm{f}\in\mathcal{F}\Big{\}},$ (9)
$\Omega_{2}(\mathcal{F})=\Big{\{}\bm{x}\rightarrow\bm{f^{\prime}}_{h_{\bm{f}}(\bm{x})}(\bm{x})~{}|~{}\bm{f}\in\mathcal{F},\bm{f^{\prime}}\in\mathcal{F}\Big{\}}$ (10)
where $\Omega_{1}(\mathcal{F})$ is the union of projections of $\mathcal{F}$ onto each dimension, and $\Omega_{2}(\mathcal{F})$ can be seen as the space of inner products from $\mathcal{F}$ and $\mathcal{F}$. Based on the above definition, the generalization bound of MUDA is given below.
@@
Let $\widehat{P}_{j}$ and $\widehat{Q}$ be the corresponding empirical distributions of ${P}_{j}$ and ${Q}$. $\widehat{\mathcal{D}}_{sj}$ and $\widehat{\mathcal{D}}_{t}$ are empirical datasets sampled from the i.i.d. samples $\mathcal{D}_{sj}=(\bm{X}_{j},\bm{Y}_{j})$ and $\mathcal{D}_{t}=\bm{X}_{t}$ of size $m$, respectively. Then, for any $\delta>0$ and $\bm{\alpha}\in\Delta=\{\bm{\alpha}:\alpha_{j}\geq 0,\sum_{j}\alpha_{j}=1\}$, with probability at least $1-3\delta$, the following holds for any scoring function $\bm{f}\in\mathcal{F}$ ,
@@
The proof of Theorem 3 can be found in Appendix -C. Theorem 3 indicates that the expected error on target domain $\mathcal{L}_{Q}\left(h_{\bm{f}}\right)$ is bounded by empirical margin error on source domain $\mathcal{L}_{{\widehat{P}}_{j}}^{\mu}$, empirical distribution gap $d^{\mu}_{\bm{f}}(\widehat{P}_{j},\widehat{S})$, the ideal error $\beta$ and rademacher complexity terms. The better generalization ability can be achieved by selecting an appropriate margin $\mu$ for margin loss. For relatively larger $\mu$ and rich hypothesis space $\mathcal{F}$, the upper bound of target error becomes tighter.

### v2

@@ Abstract
Accurate recognition of human motion intention (HMI) is beneficial for exoskeleton robots to improve the wearing comfort level and achieve natural human-robot interaction. A classifier trained on labeled source subjects (domains) performs poorly on unlabeled target subject since the difference in individual motor characteristics. The unsupervised domain adaptation (UDA) method has become an effective way to this problem. However, the labeled data are collected from multiple source subjects that might be different not only from the target subject but also from each other. The current UDA methods for HMI recognition ignore the difference between each source subject, which reduces the classification accuracy. Therefore, this paper considers the differences between source subjects and develops a novel theory and algorithm for UDA to recognize HMI, where the margin disparity discrepancy (MDD) is extended to multi-source UDA theory and a novel weight-aware-based multi-source UDA algorithm (WMDD) is proposed. The source domain weight, which can be adjusted adaptively by the MDD between each source subject and target subject, is incorporated into UDA to measure the differences between source subjects. The developed multi-source UDA theory is theoretical and the generalization error on target subject is guaranteed. The theory can be transformed into an optimization problem for UDA, successfully bridging the gap between theory and algorithm. Moreover, a lightweight network is employed to guarantee the real-time of classification and the adversarial learning between feature generator and ensemble classifiers is utilized to further improve the generalization ability. The extensive experiments verify theoretical analysis and show that WMDD outperforms previous UDA methods on HMI recognition tasks.
@@ I Introduction
The model-free methods map the sensing signals into target HMI directly without parameter identification [20, 21, 22]. The above model-free methods can achieve good performance on specific subject but fail to perform well on cross-subjects. Due to the difference in individual motor characteristics, such as kinematic properties, most current methods require labeling a large amount of data and training specific classifiers for each individual, which is burdensome [16]. Unsupervised domain adaptation (UDA) is a common method to solve the above problem [7, 6, 8, 48], where the source subjects (domains) data are labeled and the target subject (domain) data are unlabeled. Currently, the UDA methods have also been widely applied in many other fields, such as fault diagnosis [55, 63], physical information estimation [53], classification tasks [56, 57], and object identification [64, 54]. In these fields, the discrepancy between the source and target domains greatly reduces prediction performance. To alleviate the performance reduction caused by this discrepancy, single-source UDA (SUDA) utilizes the adversarial-based methods to align two domains [13, 12, 11] or explores different metric learning schemes to minimize this divergence [3, 32, 25].
In actual HMI recognition tasks, the labeled data are collected from multiple source subjects that might be different not only from the target subject but also from each other. However, the above UDA methods for HMI recognition don’t consider the difference between the source subjects, which may result in a sub-optimal solution to the problem and hinder the improvement of classifier performance [49]. Therefore, this paper considers the difference between the source subjects, and aims to propose a novel multi-source UDA (MUDA) method for HMI recognition so as to further improve the classification accuracy in the target subject. In MUDA field, there are following two key challenges that need to be addressed.
@@
Figure 1: The overview of the weighted multi-source unsupervised domain adaptation (WMDD) method for human motion intention recognition. The training data are from $N$ source subjects with labels and one target subject without labels. The features of the sensor signal are extracted by feature generator. The final goal is to use feature generator and ensemble classifiers to classify the target subject intention accurately. Firstly, the auxiliary classifiers are optimized by maximizing the discrepancy, then the source domain weight can be achieved according to the estimated MDD. Secondly, the feature generator and classifiers are trained to minimize total loss. Finally, motivated by adversarial learning, the classifiers and generator are optimized by minimizing and maximizing classifiers discrepancy.
@@
For challenge 2, the current MUDA methods [35, 33, 58, 61, 34, 62, 59, 60] introduced the source domain weight to integrate source domain differences and optimize the weighted sum loss of multiple sources to improve the accuracy of classification. The source domain weight is mainly calculated through two types: solving optimization problem about domain weight [35, 33, 58] and estimating domain weight based on the discrepancy or similarity between source and target domains [61, 34, 62, 59, 60]. However, the first type may consume more computation cost when the parameter space of optimization problem is high. The second type may not better capture the discrepancy or similarity relationship between domains [24] when source and target distributions are complex. In addition, the theoretical foundations of weight-based MUDA methods are not complete, and there is still a gap between theory and algorithms.
To solve the above problems, motivated by the margin disparity discrepancy [3] and distribution-weighted hypothesis [29], we extend MDD in SUDA that contains more generalization information to MUDA field to accurately measure the discrepancy between source and target domains, and propose a novel weighted MUDA algorithm (WMDD), where the MDD is estimated by the auxiliary classifiers and source domain weight is adjusted by the estimated MDD. Moreover, the new generalization bound of MUDA is provided and proven, and the gap between theory and algorithm is analyzed in detail. In addition, the theoretical and empirical results validate that WMDD can learn diverse features and adapt well to unknown target subject in HMI recognition tasks. Fig. 1 shows the overview of the proposed WMDD method. The contributions of this paper are given below.
@@
    A theory of MUDA is developed based on margin disparity discrepancy, and a novel weight-aware-based MUDA algorithm is proposed for HMI recognition.
@@
    A lightweight network is designed to guarantee the real-time of classification, and adversarial learning is utilized to further improve the generalization ability.
@@
II Related Work
This section presents a brief overview of the literature in the area of human motion intention (HMI) recognition, the theory of domain adaptation, multi-source unsupervised domain adaptation (MUDA), and the differences between the proposed method of this paper and the previous related works.
@@ II-A Human Motion Intention Recognition
Model-free methods based on deep networks, which can obtain higher-level features from sensor signals without domain-specific knowledge, have achieved great success in recent years, such as CNN [38], LSTM [37], CNN-BiLSTM [36]. However, the model-free methods based on the above deep networks are difficult to perform well on cross-subjects problem due to individual differences. Domain adaption is an effective method to cross-domain (subject) problems. Zhang et al.[8] incorporated an unsupervised cross-subject adaptation method to predict the HMI of the target subject without labels. Zhang et al.[7] proposed a novel non-adversarial cross-subject adaptation method (EDHKD) for HMI recognition. The ensemble diverse hypotheses method was designed to mitigate the cross-subject divergence [6]. EDHKD is the state-of-the-art method for HMI using domain adaptation technology. It mitigates the cross-subject divergence by training feature generators to minimize the upper bound of the classification discrepancy, and maximizes the discrepancy among multiple feature generators to learn diverse and complete features. However, the above methods ignore the difference between each source subject, essentially a single-source UDA method, which may cause a sub-optimal solution to the problem. Few multi-source UDA studies have focused on recognizing HMI. Therefore, this paper aims to propose a novel multi-source UDA method for HMI recognition. To the best of our knowledge, WMDD is the first work about multi-source UDA used for HMI recognition.
II-B Unsupervised Domain Adaptation Theory
@@
The study of domain adaptation theory is central of generalization error bound. They focus on providing the theoretical guarantee for prediction performance on target domain (subject). Ben-David et al.[31, 1] conducted the pioneering theoretical works in domain adaptation field. They used the $\mathcal{H}\Delta\mathcal{H}$ (the symmetric difference hypothesis space for a hypothesis space $\mathcal{H}$) divergence to replace the traditional distribution discrepancies and overcame the difficulties in estimation from finite samples. Mansour et al.[41] extended the zero-one loss of [31] to the general loss function of binary classification and developed a generalization theory. Kuroki et al.[30] proposed a more tractable source-guided discrepancy by fixing the hypothesis of [41] to the ideal source minimizer. Zhang et al.[3] extended the theory of [1] to multiple classes by introducing margin disparity discrepancy (MDD), which can characterize the difference of the multi-class scoring hypothesis. Zhang et al.[32] proposed multi-class scoring disagreement divergence based on MDD [3], that can characterize element-wise disagreements of multi-class scoring hypotheses by aggregating violations of absolute margin.
II-C Multi-source Unsupervised Domain Adaptation Algorithm
The current MUDA methods focused on aligning the distribution of each pair of source and target to reduce their domain shift by minimizing the combined discrepancy between source and target domains. The source domain weight was used to integrate source domain differences, and mainly calculated through solving related optimization problem [35, 33] or calculating the discrepancy or similarity between source and target domains [61, 34, 62, 59, 60]. For the first-type methods, Chen et al.[35] used Limited memory Broyden-Fletcher-Goldfarb-Shanno algorithm to solve domain weight, and optimized networks by minimizing the source mixture loss and the Pearson divergence. Wen et al.[33] solved domain weight through reformulating complex optimization problem, and optimized networks by minimizing source loss and discrepancy.
For the second-type methods, Yan et al.[61] calculated domain weight based on class prior distribution, and proposed weighted domain adaption network to alleviate the effect of class weight bias. Yao et al.[34] quantified the importance of different source domains and aligned the source and target distributions by minimizing maximum mean discrepancy. Wen et al.[62] proposed a maximum likelihood weight estimation approach to estimate source weight function, which matches the source of relevant part to target domain. Zuo et al.[59] solved domain weight through calculating the frequencies of target domain belonging to each source domain, and paid more attention on the source domains with higher similarities. Hu et al.[60] introduced weight allocator to map the classification loss and domain discrimination loss of each sample to its weight in weighted domain alignment loss.
However, the first-type methods would consume high computation cost when the parameter space is large since the weights are constantly optimized in training process. The above second-type methods may fail to better capture the relationship between different domains when the domain distributions are complex. The proposed method of this paper belongs to the second-type method. We determine the domain weight through the estimated MDD with less computation cost, which can be adaptively adjusted to better describe the correlation between each source domain and target domain.
II-D The Proposed Method
The proposed method of this paper is different from the above multi-source domain adaptation methods that utilize source domain weight in the below three aspects:
1) Weight determination: The source domain weight is determined based on the estimated margin disparity discrepancy, instead of class prior distribution [61], maximum likelihood estimation [62] or sample correction [59, 60]. Compared with the domain weight optimized constantly by solving unconstrained problem [35], the proposed method can reduce the computation cost and enhance algorithm stability.
2) Theoretical guarantee: There are fewer complete theoretical guarantees and analysis in the above MUDA methods [35, 33, 58, 61, 34, 62, 59, 60]. We provide the complete theoretical guarantee for MUDA based on MDD theory. The theory of WMDD is based on each single source distribution instead of the whole multi-source class-conditionals [34]. The theory of the proposed method of this paper fully considers the differences between each source domain, which is beneficial for improving classification accuracy.
3) Discrepancy measurement and real-time: The distribution disparity metric is margin disparity discrepancy that contains more generalization bound information, rather than maximum mean discrepancy [34] or Pearson divergence [35]. A lightweight network is employed to guarantee the real-time of classification, instead of using multiple complex networks [33] that is difficult to guarantee the real-time.
@@ III Theoretical Analysis of Multi-source Unsupervised Domain Adaptation
Figure 2: The comparison between the single-source unsupervised domain adaptation (SUDA) and multi-source unsupervised domain adaptation (MUDA), where different shapes represent different data categories and different colors denotes different domains. The source and target distributions of SUDA are not matched well. For MUDA, the target distribution hardly matches all subject distributions, and the discrepancy between the target distribution and each subject distribution might be different.
@@ III-A Problem Formulation
Fig. 2 compares single-source unsupervised domain adaptation (SUDA) and multi-source unsupervised domain adaptation (MUDA). In MUDA, we consider $\mathcal{C}$ class, $N$ source domains $\{\mathcal{D}_{sj}\}_{j=1}^{N}$ and one target domain $\mathcal{D}_{t}$ problem. There are two different but related distributions over $\mathcal{X}\times\mathcal{Y}$, namely source distribution $P$ and target distribution $Q$. The term $\mathcal{X}$ denotes input space and $\mathcal{Y}$ denotes output space, where $\mathcal{Y}$ is $\{0,1\}$ in binary classification.
The learner of UDA is trained on the labeled data $\mathcal{D}_{sj}=\{(\bm{x}_{i}^{sj},y_{i}^{sj})\}_{i=1}^{|\mathcal{D}_{sj}|}$ and the unlabeled data $\mathcal{D}_{t}=\{\bm{x}_{i}^{t}\}_{i=1}^{|\mathcal{D}_{t}|}$, where the labeled data are sampled from source distribution $P_{j}$ and the unlabeled data are sampled from target distribution $Q$. The goal of MUDA is to learn the hypothesis space $\mathcal{H}$ of labeling function $h:\mathcal{X}\rightarrow\mathcal{Y}$ to minimize expected target error
@@
where $L$ is the loss function. Ben-David et al.[1] used zero one loss of the form $\mathbb{I}[h(\bm{x})\neq y]$ to represent $L$, where $\mathbb{I}$ is the indicator function. Following [32], we consider the hypothesis space $\mathcal{F}$ that contains scoring function $\bm{f}:\mathcal{X}\rightarrow\mathbb{R}^{\mathcal{C}}$ which outputs the prediction confidence. Then the labeling function can be induced
@@
where $f_{y}$ denotes the $y$-th component of vector function $\bm{f}$. Zhang et al.[3] defined the margin of $\bm{f}$ for labeled sample $(x,y)$ as $\omega_{\bm{f}}(\bm{x},y)=[f_{y}(\bm{x})-\max_{y^{\prime}\neq y}f_{y^{\prime}}(\bm{x})]/2$. Then, the margin loss of $\bm{f}$ can be denoted as
@@
Margin disparity discrepancy (MDD) [3], containing more generalization information, is used to measure the distribution divergence between $P$ and $Q$, which is defined as
@@
The proof for Theorem 1 can be found in Appendix VII-A. In the following section, we give the cross-domain generalization bound for MUDA based on Theorem 1.
@@ III-B Generalization Bound for MUDA
In MUDA, some source domains, which are more relevant to the target domain than others, might be more important to classify. Therefore, we use domain weight $\alpha_{j}\geq 0$ to describe this and give the following theorem.
@@
The proof of Theorem 2 can be found in Appendix VII-B. Given the fixed $\beta$, the target error (generalization error) $\mathcal{L}_{Q}\left(h_{\bm{f}}\right)$ is determined by the distribution discrepancy $d^{\mu}_{\bm{f}}(P_{j},Q)$ and the expected loss $\mathcal{L}_{{P}_{j}}^{\mu}(\bm{f})$. In Eq. (7), the discrepancy $d^{\mu}_{\bm{f}}(P_{j},Q)$ bounds the performance gap caused by domain distribution shift. The smaller $\mathcal{L}_{{P}_{j}}^{\mu}(\bm{f})$ indicates better performance of scoring function $\bm{f}$ on the $j$-th source domain.
Note that, the above theorem only considers the ideal and complete distribution. However, the collected data fails to cover the entire distribution space in actual learning. Therefore, we further consider the sampling error and give the empirical form of Theorem 2. First, Rademacher complexity that is widely used in the generalization theory is introduced to measure the richness of a certain hypothesis space [39, 3]. The definition is given below:
@@
$\Omega_{1}(\mathcal{F})=\Big{\{}(\bm{x},y)\rightarrow\bm{f}_{y}(\bm{x})~{}\big{|}~{}\bm{f}\in\mathcal{F}\Big{\}},$ (9)
$\Omega_{2}(\mathcal{F})=\Big{\{}\bm{x}\rightarrow\bm{f^{\prime}}_{h_{\bm{f}}(\bm{x})}(\bm{x})~{}\big{|}~{}\bm{f}\in\mathcal{F},\bm{f^{\prime}}\in\mathcal{F}\Big{\}},$
where $\Omega_{1}(\mathcal{F})$ and $\Omega_{2}(\mathcal{F})$ can be regarded as the generated space based on the space $\mathcal{F}$. Then, according to the above two definitions, the generalization bound of multi-source UDA can be given in the below theorem.
@@
Let $\widehat{P}_{j}$ and $\widehat{Q}$ be the corresponding empirical distributions of ${P}_{j}$ and ${Q}$. $\widehat{\mathcal{D}}_{sj}$ and $\widehat{\mathcal{D}}_{t}$ are the empirical datasets sampled from samples $\mathcal{D}_{sj}$ and $\mathcal{D}_{t}$ with the size $m$, respectively. Then, for any $\delta>0$, any $\bm{\alpha}\in\Delta=\{\bm{\alpha}:\alpha_{j}\geq 0,\sum_{j}\alpha_{j}=1\}$ and any scoring function $\bm{f}\in\mathcal{F}$, with probability at least $1-3\delta$, the following holds,
@@
The proof of Theorem 3 can be found in Appendix VII-C. Theorem 3 indicates that the expected error on target domain $\mathcal{L}_{Q}\left(h_{\bm{f}}\right)$ is bounded by four parts, including empirical margin error on source domain $\mathcal{L}_{{\widehat{P}}_{j}}^{\mu}$, empirical distribution gap $d^{\mu}_{\bm{f}}(\widehat{P}_{j},\widehat{S})$, the ideal error $\beta$ and rademacher complexity terms. The better generalization ability can be achieved by selecting an appropriate margin $\mu$ for margin loss. Since the term $\beta$ is to take the minimum value in space $\mathcal{F}$, the upper bound of target error becomes tighter for relatively larger $\mu$ and rich hypothesis space $\mathcal{F}$. However, this bound cannot reach an acceptable small value for too large $\mu$.